OpenStudy (tifferslovesyaoi):

use the graph of y=sin2θ to find the value of sin 2θ for θ= pi/4 radians A. -1 B. 0 C. 0.5 D. 1

Nnesha (nnesha):

for theta = pi/4 substitute pi/4 for tehta \[y=\sin(2\cdot \frac{\pi}{4})\] simplify the fraction

Nnesha (nnesha):

and then graph the equation https://www.desmos.com/calculator what is y when x=pi/4

OpenStudy (tifferslovesyaoi):

1?

Nnesha (nnesha):

how would you find the period ?

OpenStudy (tifferslovesyaoi):

I honestly don't know. I am so confused by this lesson

Nnesha (nnesha):

looks like you're just guessing well here is the general equation \[\large\rm y=A \sin(Bx-C)+D\] \[\left| A \right|\] is amplitude \[P=\frac{2\pi}{B}\] 2pi because the wave repeats every 2 units (As you can see on the unit circle) and C= horizontal shift D= vertical shift

Nnesha (nnesha):

P=period use that equation to find the value of `B` when p=4pi

Nnesha (nnesha):

\[4\pi=\frac{2\pi}{B}\] solve for B (B=the number at front of x variable )

OpenStudy (tifferslovesyaoi):

I can't figure out what to put for B. I don't see a number before the x

OpenStudy (tifferslovesyaoi):

so would it be B? since there is a 4 in it?

Nnesha (nnesha):

use the equation i gave to find B

Nnesha (nnesha):

we know that Period = 2pi/b so since the given is period which is 4pi we can solve for b

Nnesha (nnesha):

\[ \color{Red}{P}=\frac{2\pi}{B}\] \[ \color{Red}{4\pi}=\frac{2\pi}{B}\] solve for B

OpenStudy (tifferslovesyaoi):

y=2*sin(2*π-4)+π y=π-2*sin(4)

Nnesha (nnesha):

no solve this equation for B \[ \color{Red}{P}=\frac{2\pi}{B}\] \[ \color{Red}{4\pi}=\frac{2\pi}{B}\] what is b = ?

Nnesha (nnesha):

given is Amplitude (which is 2) and the period (which is 4pi) we don't have to deal with C and D

OpenStudy (tifferslovesyaoi):

((4*π)/(2*π)) 2 ((2*π)/(4*π)) 1/2

Nnesha (nnesha):

that's correct B=1/2

Nnesha (nnesha):

btw in this example x is same thing as theta

Nnesha (nnesha):

so if B =1/2 how would you write the equation with A=2 and B =1/2

OpenStudy (tifferslovesyaoi):

y=0.5sin4θ

Nnesha (nnesha):

well amplitude is 2 for sure right ?? bec that is give so just replace A with 2 and we just found out that b =1/2 so replace b with 1/2

Nnesha (nnesha):

\[\huge\rm y=Asin b \theta\] without vertical and horizontal shift

OpenStudy (tifferslovesyaoi):

y=2sinθ/2?

Nnesha (nnesha):

yes

OpenStudy (tifferslovesyaoi):

thank you so much

Nnesha (nnesha):

np